# Experiment design for continuous-time systems using generalized filtering

### v1

@@ 5 Online experiment design
The basic idea to achieve (36) involves designing the input so that the rank of the data matrix increases progressively, that is, for all $k\in\{1,2,\dots,n+m-1\}$,
@@
Condition (41) can be satisfied if the input signal is designed during system operation via the following online procedure.
@@
    $\xi^{\top}\chi_{k}+\eta^{\top}\mu_{k}\neq 0.$
Then, the resulting trajectory $(u,x)\in\mathfrak{B}_{(n+m)T}$ satisfies (36).
Proposition 3 follows from [van2021beyond, Theorem 1] applied to the discrete-time system (22). Its rationale is as follows. For a piecewise constant input signal $u$, the sampled data obey the dynamics in (22). Then, by Assumption 1 and Lemma 1, at time instant $kT$, $k\in\{0,1,\dots,N-1\}$, a $\mu_{k}$ can always be designed such that (41) holds.
By combining Proposition 1 and Proposition 3, we conclude that Problem 2 is solved. We summarize this in the following theorem.
@@
Let us now go one step further. The following result shows that the online input design ensures not only that the rank condition in (36) holds, but that the stronger condition in (25) holds.
@@
According to Proposition 3, (36) holds, implying that
@@
Next, we compare Theorem 2 to Lemma 2, i.e., [lopez2022continuous, Lemma 1]. It is not difficult to see that Lemma 2 requires at least $n+m+nm$ samples to guarantee the full rank condition (25). The importance of Theorem 2 is to guarantee (25) with the minimum number of samples possible, namely, $n+m$. We further note that differently from Lemma 2, where the input signal can be designed entirely before data is collected from the system, the experiment design procedure in Proposition 3 constructs the input signal during system operation.
6 Numerical examples
In this section, we consider numerical examples of the rank condition relating sampled data to filtered data, and we demonstrate the sample efficiency of the proposed online experiment design.
Consider the aircraft longitudinal dynamics [moustakis2018adaptive] with
@@ 6 Numerical examples
$B=$ $\begin{bmatrix}0.0065&0.0534\\ -0.6103&0.0020\\ -74.6355&0.5431\\ 0&0\end{bmatrix}$,
where the four states are the velocity along the $x$- and $z$-body-axis, the angular velocity around the $y$-body-axis, and the pitch angle, while the two inputs are the elevator deflection and throttle. Note that $n+m=6$. Let $T=0.1$. Because the eigenvalues of $A$ are $\{-3.1548\pm 6.0892i,-0.0185\pm 0.3263i\}$, Assumption 1 holds. In general, although $A$ is unknown, Assumption 1 fails to hold only on a set of measure zero. Let $x(0)=\begin{bmatrix}2&-1&1&0.5\end{bmatrix}^{\top}$. By designing the input as in Proposition 3, we collect the input-state samples
@@
This is in line with the conclusion of Proposition 1. Finally, we reconstruct the system matrices as in (18)

### v2

@@ 5 Online experiment design
The basic idea to achieve $\operatorname{rank}(\bar{D})=n+m$ involves designing the input so that the rank of the data matrix increases progressively, that is, for all $k\in\{1,2,\dots,n+m-1\}$,
@@
Condition (40) can be satisfied if the input signal is designed during system operation via the following online procedure inspired by [23, Theorem 1].
@@
    $\xi^{\top}\chi_{k}+\eta^{\top}\mu_{k}\neq 0.$ (44)
Then, the resulting trajectory $(u,x)\in\mathfrak{B}_{(n+m)T}$ satisfies $\operatorname{rank}(\bar{D})=n+m$.
Proposition 3 follows from [23, Theorem 1] applied to the discrete-time system (22). Its rationale is as follows. For a piecewise constant input signal $u$, the sampled data obey the dynamics in (22). Then, by Assumption 1 and Lemma 1, at time instant kTkT, $k\in\{0,1,\dots,N-1\}$, a $\mu_{k}$ can always be designed such that (40) holds. Note that the selection of $\mu_{k}$ in (44) is in general not unique: one may exploit such freedom to select a solution which satisfies a norm constraint, e.g., to account for actuation constraints.
We are now in a position to solve Problem 2 using the following theorem.
@@
Proof. Since $u$ is designed according to Proposition 3, the resulting trajectory $(u,x)\in\mathfrak{B}_{(n+m)T}$ satisfies $\operatorname{rank}(\bar{D})=n+m$. Then, we conclude from Proposition 1 that (17) is satisfied since $N=n+m$ and $\operatorname{rank}(\bar{F})=n+m$. $\Box$
Let us now go one step further. The following result shows that the online input design ensures not only $\operatorname{rank}(\bar{D})=n+m$, but that the stronger condition in (25) holds.
@@
According to Proposition 3, we have $\operatorname{rank}(\bar{D})=n+m$, implying that
@@
Next, we compare Theorem 2 to Lemma 2, i.e., [10, Lemma 1]. It is not difficult to see that Lemma 2 requires at least $n+m+nm$ samples to guarantee the full rank condition (25). The importance of Theorem 2 is to guarantee (25) with the minimum number of samples possible, namely, $n+m$. We further note that differently from Lemma 2, where the input signal can be designed entirely before data is collected from the system, the experiment design procedure in Proposition 3 constructs the input signal during system operation.
6 Numerical example
In this section, we consider a numerical example to demonstrate that, using the experiment design method in Section 5, both the sampled data and the filtered data generated by different filters are informative for system identification. The example is concluded by showing that the system matrices can be reconstructed from the filtered data.
The system to be identified is described by the aircraft longitudinal dynamics in [11] with system matrices
@@ 6 Numerical example
$B=$ $\begin{bmatrix}0.0065&0.0534\\ -0.6103&0.0020\\ -74.6355&0.5431\\ 0&0\end{bmatrix}$.
The four states of the system are the velocity along the $x$- and $z$-body-axis, the angular velocity around the $y$-body-axis, and the pitch angle, while the two inputs are the elevator deflection and throttle. Note that $n+m=6$. Let $T=0.1$. Because the eigenvalues of AA are $\{-3.1548\pm 6.0892i,-0.0185\pm 0.3263i\}$, Assumption 1 holds. In general, even though AA is unknown, Assumption 1 fails to hold only on a set of measure zero. Let $x(0)=\begin{bmatrix}2&-1&1&0.5\end{bmatrix}^{\top}$. By designing the input as in Proposition 3, we collect the input-state samples
@@
This is in line with the conclusion of Proposition 1. Being the filtered data informative for system identification, we can finally reconstruct the system matrices as in (18):